# Quantum Statistical Effects on Warm Dark Matter and the Mass Constraint from the Cosmic Large Scale Structure

### v1

@@ A.2 The background evolution
The two independent dynamical variables are $x$ and $\Delta$ ($x$ and $r$ for the boson case with $\Delta=0$), whose background evolution is derived using the conservation of the comoving particle number and the conservation of the specific entropy (concerning the total particle number), that is
@@
The initial conditions are described as the following. For both the fermion and boson cases, the value of $x/a$ is a constant in the early universe. Thus, we specify the value of $x/a$ in the relativistic limit as the first initial condition. We call $T_{0}/m\equiv\lim_{a\rightarrow 0}a/x$ the initial “warmness” of qWDM. The second initial condition is treated differently for two possibilities. For the fermion case and the boson case without a BEC component, the value of $\Delta$ is a constant in the early universe. In this case, we specify $\Delta$ in the relativistic limit ($\Delta_{\rm i}$) as the second initial condition. For the boson case with $\Delta=0$, a BEC component appears. The BEC fraction $r$ is constant in the early universe. In that case, we use $r$ in the relativistic limit ($r_{\rm i}$) as the second initial condition. That is, the initial conditions are specified by the set of $(x_{\rm i}/a_{\rm i},\,\Delta_{\rm i})$, except for the boson case with $\Delta_{\rm i}=0$ where they are specified by $(x_{\rm i}/a_{\rm i},\,r_{\rm i})$.
With the background evolution of $x$ and $\Delta$ obtained above, we substitute them into Eqs. (A4)-(A6) to get the background evolution of the number density, energy density and pressure of qWDM. We assume the qWDM perturbation is an adiabatic process with the adiabatic sound speed calculated by Eq. (10). With a finite equation of state and a finite adiabatic sound speed, the qWDM linear perturbation equations are modified compared to the cold DM case and read (Ma & Bertschinger, 1995),
@@
where $\delta_{{\rm q}{\textsc{wdm}}}$ and $q_{{\rm q}{\textsc{wdm}}}$ are the qWDM overdensity and heat flux, denotes the derivative with respect to the conformal time, $w_{{\rm q}{\textsc{wdm}}}\equiv\frac{P}{\rho_{\rm tot}}$, $\mathcal{H}$ is the conformal Hubble parameter, and $h_{\rm s}$ and $\eta_{\rm s}$ are the two synchronous gauge gravitational variables (Ma & Bertschinger, 1995). We have omitted the pressure anisotropy in Eq. (A15) since we assume qWDM has an equilibrium phase-space distribution that is isotropic. We implement the above linear perturbation into the public code camb (Lewis et al., 2000) to calculate the linear matter power spectrum and determine the suppression scale using Eq. (18).

### v2

@@ A.2 The background evolution
The two independent dynamical variables are $x$ and $\Delta$ ($x$ and $r$ for the boson case with $\Delta=0$), whose background evolution is derived using the conservation of the comoving particle number and the conservation of the specific entropy (concerning the total particle number), that is,
@@
The initial conditions are described as follows. For both the fermion and boson cases, the value of $x/a$ is a constant in the early Universe. Thus, we specify the value of $x/a$ in the relativistic limit as the first initial condition. We call $T_{0}/m\equiv\lim_{a\rightarrow 0}a/x$ the initial “warmness” of qWDM. The second initial condition is treated differently for two possibilities. For the fermion case and the boson case without a BEC component, the value of $\Delta$ is a constant in the early Universe. In this case, we specify $\Delta$ in the relativistic limit ($\Delta_{\rm i}$) as the second initial condition. For the boson case with $\Delta=0$, a BEC component appears. The BEC fraction $r$ is constant in the early Universe. In that case, we use $r$ in the relativistic limit ($r_{\rm i}$) as the second initial condition. That is, the initial conditions are specified by the set of $(x_{\rm i}/a_{\rm i},\,\Delta_{\rm i})$, except for the boson case with $\Delta_{\rm i}=0$ where they are specified by $(x_{\rm i}/a_{\rm i},\,r_{\rm i})$.
With the background evolution of $x$ and $\Delta$ obtained above, we substitute them into Eqs.,(A4)-(A6) to get the background evolution of the number density, energy density, and pressure of qWDM. We assume that the qWDM perturbation is an adiabatic process with the adiabatic sound speed calculated by Eq.,(11). With a finite equation of state and a finite adiabatic sound speed, the qWDM linear perturbation equations are modified compared to the cold DM case and read (Ma & Bertschinger, 1995),
@@
where $\delta_{{\rm q}{\textsc{wdm}}}$ and $q_{{\rm q}{\textsc{wdm}}}$ are the qWDM overdensity and heat flux, denotes the derivative with respect to the conformal time, $w_{{\rm q}{\textsc{wdm}}}\equiv\frac{P}{\rho_{\rm tot}}$, $\mathcal{H}$ is the conformal Hubble parameter, and $h_{\rm s}$ and $\eta_{\rm s}$ are the two synchronous gauge gravitational variables (Ma & Bertschinger, 1995). We have omitted the pressure anisotropy in Eq.,(A15) since we assume that qWDM has an equilibrium phase-space distribution that is isotropic. We implement the above linear perturbation into the public code camb (Lewis et al., 2000) to calculate the linear matter power spectrum and determine the suppression scale using Eq. (19).
Appendix B The self-interaction strength
We have investigated the effects of quantum statistics, including the shift of time of RNRT and the corresponding impacts in the matter power spectrum. For those effects to hold, a certain level of self-interaction is required to maintain DM equilibrium until RNRT is fully completed. Specifically, DM must remain in equilibrium until a scale factor $a_{\rm f}$ sufficiently later than the time of RNRT, such as $a_{\rm f}>10a_{\rm nr}$. The necessary self-interaction strength depends on the type of interaction. In this context, we consider a DM self-interaction with a constant cross section $\sigma_{\textsc{si}}$ to demonstrate that this requirement can be met without violating current astrophysical and cosmological constraints.
Figure 6: The minimum constant cross section required for $a_{\rm f}>10a_{\rm nr}$ for the fermion case (left) and the boson case with BEC (right). Different curves correspond to different degrees of degeneracy for the fermion cases and different BEC fraction for the boson cases. The dots on curves mark the corresponding mass low-bound from the Lyman-$\alpha$ forest for different cases. The required minimum cross section is well below the astrophysical upper bounds ranging $0.2\sim 0.5$ cm2/g (Randall et al., 2008; Eckert et al., 2022). Considering a velocity-dependent cross section with $\sigma_{\textsc{si}}\propto\langle v\rangle^{-p}$ for some positive $p$ raises the astrophysical constraints on the cross section in the near-relativistic regime.
We estimate $a_{\rm f}$ by equaling the DM self-interaction rate per particle in thermal states with the Hubble rate, i.e., $\sigma_{\textsc{si}}\langle v\rangle n_{\rm tot}\big{|}_{a_{\rm f}}=H\big{|}_{a_{\rm f}}$. Like the sound speed, we parameterized average velocity at the nonrelativistic phase as $\langle v\rangle=\frac{a_{\rm nr}}{\sqrt{3}a}$. With $a_{\rm nr}$ solved in Eq. (16), it can be shown that
$a_{\rm f}=5.6\times 10^{-4}\left(\frac{2/g}{\gamma(\Delta_{\rm i},r_{\rm i})}\right)^{\nicefrac{{1}}{{6}}}\left(\frac{\Omega_{\rm dm}h^{2}}{0.12}\right)^{\nicefrac{{2}}{{3}}}\left(\frac{\sigma_{\textsc{si}}/m}{1\,{\rm{cm^{2}/g}}}\right)^{\nicefrac{{1}}{{2}}}\left(\frac{1\,\rm{keV}}{m}\right)^{\nicefrac{{2}}{{3}}}\,.$ (B1)
The above result is consistent with that given in Garani et al. (2022) in the fiducial case where $\Delta_{\rm i}=0$ and $r_{\rm i}=0$ but is generalized to include scenarios involving degenerate fermions and bosons with BEC. The minimally required interaction strength is shown in Figure 6, presented by $\sigma_{\textsc{si}}/m$ as a function of qWDM mass. The minimally required cross section varies with the degree of degeneracy for the fermion case and the BEC fraction for the boson case. Considering the mass constraints from the Lyman-$\alpha$ forest, the minimally required $\sigma_{\textsc{si}}/m$ for the lowest allowed mass is about $1.4\times 10^{-6}~{}{\rm cm}^{2}/{\rm g}$ (see the height of the dots in Figure 6) and is lower for a higher DM particle mass. These are well below the upper bound from the current astrophysical constraints $0.2\sim 0.5~{}{\rm cm}^{2}/{\rm g}$ (Randall et al., 2008; Eckert et al., 2022).
In the above discussion, we have assumed a velocity-independent cross section. However, the conclusion that the required cross section well satisfies the current astrophysical constraints still holds as long as $\sigma_{\textsc{si}}\propto\langle v\rangle^{-p}$ with some positive $p$ for the following reason. Since we require DM remain in equilibrium until RNRT is complete, the above required cross section should be interpreted as that for relativistic or nearly relativistic case. The current astrophysical constraints are at $v\sim 1000$ km/s for the case in Randall et al. (2008) and at $v\sim 5000$ km/s for the case in Eckert et al. (2022). These will be translated into a higher upper bound of $\sigma_{\textsc{si}}/m$ in the nearly-relativistic case. Compared to current astrophysical constraints, only a weak self-interaction is necessary to maintain the equilibrium of qWDM until RNRT is complete. This is primarily due to the high number density of qWDM in the early Universe, which results in a significant self-interaction rate, even with a small cross section. Consequently, the strong self-interactions often assumed for WDM would keep it in its own thermal equilibrium, preventing it from behaving as a free-streaming particle.